OpenStudy (anonymous):

so if I have an equation with two fractions in it and two whole numbers and one of the fractions has an x next to it in order for me to get rid of those fractions I multiply by the LCD but do I also multiply the whole numbers by that?

jimthompson5910 (jim_thompson5910):

what is the equation

OpenStudy (anonymous):

well like the one we just did. what if there was a whole number in there would I have to also times that by the LCD?

jimthompson5910 (jim_thompson5910):

yes you need to multiply every term by the LCD so things balance out

jimthompson5910 (jim_thompson5910):

but what is the equation?

OpenStudy (anonymous):

2x+3=3/4x--2

OpenStudy (phi):

\[ 2x+3=\frac{3}{4}x - -2 \] I would immediately write -(-2) as +2 to avoid mistakes \[ 2x+3=\frac{3}{4}x +2 \] to get rid of the 4, multiply both sides of the equation (and *all* terms) by 4. like this \[ 4\cdot 2x+ 4\cdot 3= 4\cdot \frac{3}{4}x + 4\cdot 2 \] now simplify

OpenStudy (phi):

the first term is 4*2*x which is 8*x , usually written 8x 4*3 is 12 \[ 4 \cdot \frac{3}{4} \cdot x = \frac{4\cdot 3\cdot x}{4} \text{ or } \frac{4}{4}\cdot \frac{3x}{1}= 3x\]

OpenStudy (phi):

the last term 4*2 is 8 you get 8x +12 = 3x +8